Jojoknph:

1) Pick a whole number 2) Double your number. 3) Add 20 to the above result. 4) Multiply the above by 5 5) Subtract 100 6) Divide by 10 The result is always the number that you started with! Suppose you start by picking N. Which of the equations below best demonstrates that the result after Step 6 is also N? A N∗2+20∗5−100÷10=N B ((2∗N+20)∗5−100)÷10=N C (N+N+20)∗5−100÷10=N D (((N÷10)−100)∗5+20)∗2=N

Laylalyssa:

is this a trick question

Jojoknph:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa is this a trick question \(\color{#0cbb34}{\text{End of Quote}}\) It shouldn't be

jhonyy9:

1) Pick a whole number 2) Double your number. 3) Add 20 to the above result. 4) Multiply the above by 5 5) Subtract 100 6) Divide by 10 x 2x 2x+20 (2x+20)*5 = 10x +100 10x +100 -100 = 10x 10x/10 = x x proven - ty.

jhonyy9:

@Laylalyssa opinion pls. ? ty

Laylalyssa:

1) Pick a whole number 5 2) Double your number. 10 3) Add 20 to the above result. 30 4) Multiply the above by 5 150 5) Subtract 100 50 6) Divide by 10 5 Basically its the same thing

jhonyy9:

yes but with N or x is in more general case